eliminate the arbitrary constant using any method

13.x^2y^3\:+\:x^3y^5\:=\:C


1
Expert's answer
2021-09-06T16:07:16-0400

"x^2y^3+x^3y^5=c"

Differentiating both sides w.r.t "x" , we get

"\frac{d}{dx}(x^2y^3+x^3y^5)=\frac{d}{dx}(c)
\\\Rightarrow \frac{d}{dx}(x^2y^3)+\frac{d}{dx}(x^3y^5)=0
\\\Rightarrow 2xy^3+x^2.3y^2\frac{dy}{dx}+3x^2y^5+x^3.5y^4\frac{dy}{dx}=0
\\\Rightarrow \frac{dy}{dx}(3x^2y^2+5x^3y^4)+2xy^3+3x^2y^5=0
\\\Rightarrow \frac{dy}{dx}=-\frac{2xy^3+3x^2y^5}{3x^2y^2+5x^3y^4}"